# Integrated Utilization of Equations and Small Dataset in the Koopman Operator: Applications to Forward and Inverse Problems

### v1

@@ II Brief Review of Previous Works
II.1 Koopman operator
II.1.1 Snapshot pairs
@@
II.1.2 Koopman operator, dictionary, and EDMD
@@
In Eq. (4), the left-hand side represents the action of the Koopman operator $\mathcal{K}$ on the observable function $\phi$, while the right-hand side represents the composition of the observable function $\phi$ with the time evolution function $F$, illustrating the evolution of observable functions in the infinite-dimensional space. Equation (5) means the following fact: due to the time-evolution of the observable function, one can evaluate the value of the observable function at the coordinates $\bm{y}_{i}$ after time evolution, $\phi(\bm{y}_{i})$, by substituting the coordinates $\bm{x}_{i}$ for the function $\mathcal{K}\phi$.
@@
$\bm{\psi}(\bm{x})=\begin{bmatrix}\psi_{0}(\bm{x})&\psi_{1}(\bm{x})&\cdots&\psi_{N_{\mathrm{dic}}}(\bm{x})\end{bmatrix}^{\top},$ (6)
@@
where $\{a_{n}^{\phi}\}_{n=1}^{N_{\mathrm{dic}}}$ represents the coefficients for the expansion of the observable function in terms of the dictionary functions. Although Eq. (7) could be an approximation rather than equality when the number of dictionary functions is insufficient, we assume, for simplicity, that equality holds in the following discussions. Applying the Koopman operator $\mathcal{K}$ to the observable function $\phi(\bm{x})$, we have
@@ III.4.1 Duffing equation
Figure 3: (Color online) Prediction errors for the Duffing equation. The number of data is 10, and the prediction error was evaluated. We performed 100 trials. The solid line corresponds to the mean of the prediction errors of the proposed method, and the shaded area on the lower side represents the quartile range. The shaded area on the upper side represents the quartile range derived from the results of the conventional EDMD algorithm; since the mean of the prediction errors show divergent behavior immediately, we only draw up to time step 3. We did not draw the dashed line representing the mean of the prediction errors because it goes beyond this drawing range and shows divergent behavior.
We repeated the above prediction error estimation 100 times using different training datasets. Figure 3 shows the prediction errors; the solid line corresponds to the mean of the prediction errors of the 100 trials, and the shaded area on the lower side represents the quartile range between the 25th and 75th percentiles. For comparison, the results of the conventional EDMD algorithm using $10$ data are also depicted; the mean of the prediction errors show divergent behavior immediately, and only the quartile range is depicted. The relationship between the interquartile range and the mean of the prediction errors indicates that the prediction results of some trials are very poor, and we confirmed that the prediction was unstable in regions away from the training dataset. In contrast, the proposed method shows a small error with a narrow range. The reason is that the prior Koopman matrix contains the information for the time-evolution, even though the parameter values are different from the true ones.
@@ III.4.2 Van der Pol equation
Figure 5: (Color online) Prediction errors for the van der Pol equation. The number of data is 10. The solid line corresponds to the mean of the prediction errors of the proposed method, and the dashed line is the result of the conventional EDMD algorithm. The shaded areas correspond to the interquartile ranges. As in Fig. 3, the conventional EDMD yielded divergent behavior, and then, we only draw up to time step 3.
@@ IV.1.2 Numerical results for the van der Pol equation
From Fig. 8, we see that the fitting results calculated for the various parameters $\{\epsilon_{n}\}$ intersect at a single point, i.e., $\hat{\mu}=0.999$. This intersection point corresponds to the estimated value, which is in close agreement with the true value $\mu_{\mathrm{true}}=1.0$. In addition, this intersection point has the mean error value of almost zero. From this reason, we can say that this intersection point gives the estimate.
Why does the intersection point yield the estimated parameter value? Of course, one of the reasons is that the point with the zero mean error should be common. There would be another reason, as follows. If we use the true parameter value as the initial one in the proposed method, the prior knowledge exactly corresponds to the dataset. Hence, the update procedure will have essentially no effect. Since the parameter $\epsilon$ affects the update, the intersection point means the update has no effect. Therefore, the intersection yields an estimation for the parameter.
@@ Appendix A Evaluate Koopman matrix from equations
In this work, we consider the monomial functions with maximum degree $5$. Hence, the number of the coupled ordinary differential equations is $21$. Solving the $21$ coupled ordinary differential equations with a specific initial condition in Eq. (34) yields a column of the Koopman matrix. Repeating this process for all initial conditions constructs the full Koopman matrix.

### v2

@@ II Brief Review of Previous Works
II.1 Snapshot pairs
@@
II.2 Koopman operator, dictionary, and EDMD
@@
In Eq. (4), the left-hand side represents the action of the Koopman operator $\mathcal{K}$ on the observable function $\phi$, while the right-hand side represents the composition of the observable function $\phi$ with the time-evolution function $F$, illustrating the evolution of observable functions in the infinite-dimensional space. Equation (5) means the following fact: due to the time-evolution of the observable function, one can evaluate the value of the observable function at the coordinates $\bm{y}_{i}$ after time-evolution, $\phi(\bm{y}_{i})$, by substituting the coordinates $\bm{x}_{i}$ for the function $\mathcal{K}\phi$.
@@
$\bm{\psi}(\bm{x})=\begin{bmatrix}\psi_{1}(\bm{x})&\psi_{2}(\bm{x})&\cdots&\psi_{N_{\mathrm{dic}}}(\bm{x})\end{bmatrix}^{\top},$ (6)
@@
where $\{a_{n}^{\phi}\}_{n=1}^{N_{\mathrm{dic}}}$ represents the coefficients for the expansion of the observable function in terms of the dictionary functions. Although Eq. (7) could be an approximation rather than an equality when the number of dictionary functions is insufficient, we assume, for simplicity, that the equality holds in the following discussions. Applying the Koopman operator $\mathcal{K}$ to the observable function $\phi(\bm{x})$, we have
@@ III.4.1 Duffing equation
Figure 3: (Color online) Prediction errors for the Duffing equation. The number of data is 10, and the prediction error was evaluated. We performed 100 trials. The solid line corresponds to the mean of the prediction errors of the proposed method, and the shaded area on the lower side represents the quartile range. The shaded area on the upper side represents the quartile range derived from the results of the conventional EDMD algorithm; since the mean of the prediction errors show divergent behavior immediately, we only draw it up to time step 3. We did not draw the dashed line representing the mean of the prediction errors because it goes beyond this drawing range and shows divergent behavior.
We repeated the above prediction error estimation 100 times using different training datasets. Figure 3 shows the prediction errors; the solid line corresponds to the mean of the prediction errors of the 100 trials, and the shaded area on the lower side represents the quartile range between the 25th and 75th percentiles. For comparison, the results of the conventional EDMD algorithm using $10$ data are also depicted; the mean of the prediction errors shows divergent behavior immediately, and only the quartile range is depicted. The relationship between the interquartile range and the mean of the prediction errors indicates that the prediction results of some trials are very poor, and we confirmed that the prediction was unstable in regions away from the training dataset. In contrast, the proposed method shows a small error with a narrow range. The reason is that the prior Koopman matrix contains the information for the time-evolution, even though the parameter values are different from the true ones.
@@ III.4.2 Van der Pol equation
Figure 5: (Color online) Prediction errors for the van der Pol equation. The number of data is 10. The solid line corresponds to the mean of the prediction errors of the proposed method, and the dashed line is the result of the conventional EDMD algorithm. The shaded areas correspond to the interquartile ranges. As in Fig. 3, the conventional EDMD yielded divergent behavior, and then, we only draw it up to time step 3.
@@ IV.1.2 Numerical results for the van der Pol equation
From Fig. 8, we see that the fitting results calculated for the various parameters $\{\epsilon_{n}\}$ intersect at a single point, i.e., $\hat{\mu}=0.999$. This intersection point corresponds to the estimated value, which is in close agreement with the true value $\mu_{\mathrm{true}}=1.0$. In addition, this intersection point has the prediction error value of almost zero. From this reason, we can say that this intersection point gives the estimate.
Why does the intersection point yield the estimated parameter value? Of course, one of the reasons is that the point with the zero prediction error should be common. There would be another reason, as follows. If we use the true parameter value as the initial one in the proposed method, the prior knowledge exactly corresponds to the dataset. Hence, the update procedure will have essentially no effect. Since the parameter $\epsilon$ affects the update, the intersection point means the update has no effect. Therefore, the intersection yields an estimation for the parameter.
@@ Appendix A Evaluate Koopman matrix from equations
In this work, we consider the monomial functions with maximum degree $5$. Hence, the number of the coupled ordinary differential equations is $21$. Solving the $21$ coupled ordinary differential equations with a specific initial condition in Eq. (34) yields a column of the Koopman matrix. Repeating this procedure for all initial conditions constructs the full Koopman matrix.